# Shamier Settle, Matthew Shapiro, Emerita Torres, and Anand Vidyashankar for discussion and feedback on the work described

### v2

@@ 4. We examine the validity of our assumptions using additional information
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is consistent. A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. This assumption is necessary to derive the standard errors.
@@
In this section, we examine each assumption. Since neither can be evaluated from the survey data alone, we consider additional information. In Section 4.1, we compare the spatial distribution of the respondents to data from the American Community Survey Public Use Micro Sample 2017-2021 and the New York City Office of Administrative Trials and Hearings. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets. [^6] The results of this section can also be derived from a spatial Negative Binomial Lévy Process in a manner that mirrors Section 3.2. The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
4.1 The Street Vendor Project survey is spatially representative
@@
If this assumption is violated, then $\hat{\theta}$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}$ would likely underestimate $\theta$ even in large samples.
We think this assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid. We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
Nevertheless, it is possible that canvassing operations systematically missed locations, and these missed locations may be revealed by examining other data sources. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data publicly available from the New York City Office of Administrative Trials and Hearings (OATH). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the five year American Community Survey Public Use Micro Sample, 2017-2021 (ACS).
@@
We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated. That is because this undercount may be compensated by an overcount in other areas (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
4.2 We model the arrival of street vendors as a pure birth process
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because the consistency of $\hat{\theta}$ holds under the wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated. [^7] The ratio estimator may no longer be asymptotically efficient as well. i.e., In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies the ratio estimator is asymptotically efficient. However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
@@ 4.2 We model the arrival of street vendors as a pure birth process
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925). That is, let $N^{t}_{i}(A)$ denote the number of vendors in area $A$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(A)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(A)$, would follow a negative binomial distribution with mean $\mu_{i}=\exp(\lambda_{i})\,M(A)$ and variance $\mu_{i}\bigr{(}\mu_{i}-M(A)\bigr{)}\,/\,M(A)$. i.e.,
$N_{i}(A)\sim\ \text{Negative Binomial}\left(\mu_{i},\ \mu_{i}\,\frac{\mu_{i}-M(A)}{M(A)}\right)$
where we parameterize the negative binomial distribution by its mean and variance. In this case, $N_{i}(A)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details. [^9] Note that we define the negative binomial and negative hypergeometric distributions as the total number of trials required to get a predetermined number of successes. The negative binomial assumes sampling with replacement, while the negative hypergeometric distribution assumes sampling without replacement.
@@
$n_{i}(A)\sim\ \text{Negative Binomial}\left(p\,\mu_{i},\ p\,\mu_{i}\,\frac{\mu_{i}-M(A)}{M(A)}\right)$
The conditional distribution of $n_{1}(A)$ given $N_{1}(A)$ is then
$n_{1}(A)\ |\ N_{1}(A)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(A)-1,\ M(A)-1,\ p\,M(A)\bigl{)}$
@@
$\mathbb{E}\left[n_{1}(A)\ |\ N_{1}(A)\right]=p\,N_{1}(A)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(A)\ |\ N_{1}(A)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(A)$
so that under increasing domain asymptotics, the standard error of
$\hat{\theta}=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
is asymptotically normal with mean $\theta=\mu_{0}$ and standard error
$\text{SE}(\hat{\theta})=\theta\sqrt{w_{0}\left(\frac{1}{p\,\theta}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
@@
$w_{0}=\dfrac{\theta-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.3 for details.
Note that this standard error is identical to the standard error from Section 3.2 except for the weights $w_{0}$ and $w_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—i.e., the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
The weights $w_{0}$ and $w_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\theta$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most be 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market only contains one or two vendors.
@@ 5. Discussion
We estimate 23,000 vendors work in New York City. In comparison, the 2017-2021 ACS estimates 4,634 people work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City. This suggests that according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicate that the ACS estimate misses 80% of street vendors or more.
A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 488, which we obtained from the replicate weights. Using the lower limit of a 95% confidence interval for our estimate (19,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly two thirds of street vendors or more.
Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population. For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 225 vendors in the average market. We believe this is highly unlikely, suggesting that even if the Poisson assumption is violated, the two estimates are inconsistent.
Additional evidence comes from the OATH data. The OATH data indicate that approximately 2,500 (unique) vendors are summoned to court for allegedly violating laws related to vending each year. If the ACS estimate of at most 4,634 vendors holds, then more than half of all vendors are summoned to court each year. If our estimate of 23,000 vendors holds, then approximately one in ten vendors are summoned to court each year. We believe the latter is more realistic.
@@
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations. For example, the ACS estimate is the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Indeed, street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy. In this case, either estimate may be relevant, depending on the intended use case.

### v4

@@ 4. Model Validation
We make two assumptions in Section 3. The main assumption is that the status and spatial distribution of the respondents is representative. This assumption is necessary to ensure the ratio estimator is accurate in large samples (i.e., consistent). A secondary assumption is that the spatial distribution of vendors is well described by a family of inhomogeneous spatial Poisson processes. This assumption is necessary to derive the standard errors.
@@
In this section, we examine each assumption. Since neither can be evaluated from the survey data alone, we consider additional information. In Section 4.1, we compare the spatial distribution of the respondents to data from the New York City Office of Administrative Trials and Hearings and the American Community Survey. In Section 4.2, we model the arrival of street vendors as a pure birth process, and we derive a more general formula for the standard errors in which the amount of extra-Poisson variation can be determined by the number of vendors that cluster within markets. [^5] The results of this section can also be derived from a spatial Negative Binomial Lévy process in a manner that mirrors Section 3.2. The advantage of the birth process approach taken in Section 4.2 is that it provides a plausible explanation for why the Poisson assumption made in Section 3.2 may not hold in practice.
4.1 Representativeness
@@
If this assumption is violated, then $\hat{\theta}(A)$ may be inconsistent. For example, if respondents are more likely to come from locations in which a higher proportion of vendors have permits, then the ratio $n_{0}(A)\,/\,n_{1}(A)$ would not be representative, and $\hat{\theta}(A)$ would likely underestimate $\theta(A)$ even in large samples.
We think the representativeness assumption is reasonable because the Street Vendor Project canvassed New York City to distribute a large amount of relief aid. We argue that trust in SVP and its mission, along with the aid, removed many barriers that typically prevent survey operations from reaching hard-to-reach populations. Instead, nonresponse reflects chance variation in the personal circumstances of a vendor that are largely unrelated to location.
Nevertheless, it is possible that canvassing operations systematically missed locations, and these missed locations may be revealed by examining other data sources. We thus compare the spatial distribution of vending locations reported by respondents to the spatial distribution of vendors who allegedly violated street vending laws during the year 2021. We consider the location of an individual’s first violation in 2021, reported in administrative data from the NYC Office of Administrative Trials and Hearings (OATH) (accessed 2024-04-03). We also compare the residencies reported by the respondents to the residencies of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers (Standard Occupational Classification 41-9091) estimated in the 2017-2021 American Community Survey (ACS) Public Use Micro Sample from the U.S. Census Bureau (accessed 2022-12-22).
@@
We also note that even if this discrepancy suggests a location was missed by the Street Vendor Project, our main assumption is not necessarily violated. That is because the undercount in some areas may be compensated by an overcount in others (e.g., the Upper East Side of Manhattan or East Brooklyn) such that our overall estimate is still accurate.
4.2 Overdispersion
The second assumption in Section 3.2 is that the spatial distribution of vendors is well approximated by a family of inhomogeneous Poisson processes. We consider this assumption secondary because the consistency of $\hat{\theta}(A)$ holds under a wide class of point processes described in Section 4.1. That said, the standard errors derived in Section 3.2 do not necessarily hold when the Poisson assumption is violated. [^6] The ratio estimator may no longer be asymptotically efficient as well. In Section 7.3, we show that under the Poisson process assumption, the ratio estimator maximizes the likelihood, which implies it is asymptotically efficient. However, the ratio estimator may not maximize the likelihood when the Poisson process assumption is violated.
@@ 4.2 Overdispersion
Now suppose instead that larger markets grow faster than smaller markets, an empirical phenomenon known as preferential attachment. Then the time between vendor arrivals may be better approximated by an exponential distribution with a linear arrival rate, or Yule process (Yule 1925). That is, let $N^{t}_{i}(B)$ denote the number of vendors in subregion $B$ at time $t$ so that the time between vendor arrivals at each market follows an exponential distribution with rate $\lambda_{i}\,N^{t}_{i}(B)$. If we assume further that each market started with one vendor at $t=0$, then the number of vendors at the time the survey was conducted ($t=1$), $N_{i}(B)$, would follow a negative binomial distribution with mean $\mu_{i}(B)=\exp(\lambda_{i})\,M(B)$ and variance $\mu_{i}(B)\bigr{(}\mu_{i}(B)-M(B)\bigr{)}\,/\,M(B)$. i.e.,
$N_{i}(B)\sim\ \text{Negative Binomial}\left(\mu_{i}(B),\ \mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
where we parameterize the negative binomial distribution by its mean and variance. In this case, $N_{i}(B)$ exhibits extra-Poisson variation, and the results of Section 3.2 do not hold. See Ross (2014, chap. 6) for details. [^7] Note that we define the negative binomial and negative hypergeometric distributions as the total number of trials required to get a predetermined number of successes. The negative binomial assumes sampling with replacement, while the negative hypergeometric distribution assumes sampling without replacement.
@@
$n_{i}(B)\sim\ \text{Negative Binomial}\left(p\,\mu_{i}(B),\ p\,\mu_{i}(B)\,\frac{\mu_{i}(B)-M(B)}{M(B)}\right)$
The conditional distribution of $n_{1}(B)$ given $N_{1}(B)$ is then
$n_{1}(B)\ |\ N_{1}(B)\sim\ \text{Negative Hypergeometric}\bigl{(}N_{1}(B)-1,\ M(B)-1,\ p\,M(B)\bigl{)}$
@@
$\mathbb{E}\left[n_{1}(B)\ |\ N_{1}(B)\right]=p\,N_{1}(B)\qquad\text{and}\qquad\mathbb{V}\bigr{(}n_{1}(B)\ |\ N_{1}(B)\bigr{)}=w_{1}\,p\,(1-p)\,N_{1}(B)$
so that under increasing domain asymptotics, the ratio estimator
$\hat{\theta}(A)=\frac{N_{1}(A)\ n_{0}(A)}{n_{1}(A)}$
is asymptotically normal with mean $\theta(A)=\mu_{0}(A)$ and standard error
$\text{SE}[\hat{\theta}(A)]=\theta(A)\sqrt{w_{0}\left(\frac{1}{p\,\theta(A)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
@@
$w_{0}=\dfrac{\theta(A)-M(A)}{M(A)}\qquad\text{and}\qquad w_{1}=\dfrac{N_{1}(A)-M(A)}{M(A)+1}$
See Section 7.3.4 for details. Note that the weights $w_{0}$ and $w_{1}$ depend on $A$, although this dependence is not reflected in the notation.
The standard error derived in this section for $\hat{\theta}(A)$ is identical to the standard error for $\hat{\theta}(A)$ from Section 3.2 except for the weights $w_{0}$ and $w_{1}$. These weights are the result of the extra-Poisson variation induced by preferential attachment—that is, the linear “birth” rate in which the arrival of new vendors is proportional to the number of vendors in the market.
The weights $w_{0}$ and $w_{1}$ correspond to the relative number of vendors without permits and with permits per market respectively. These weights cannot be estimated from the survey data alone. However, we believe that $\theta(A)$ and $N_{1}(A)$ are no more than 10 times larger than $M(A)$. It follows that $w_{0}$ and $w_{1}$ are less than 9, and thus the standard errors from Section 3.2 and the margins of error from Table 1 are at most 3 times larger if the Poisson assumption is violated and the Yule process holds. We suspect this bound is conservative, however, and the typical market contains only one or two vendors.
The same behavior is exhibited by the subregion estimator of $\theta(B)$,
$\tilde{\theta}(B)=\dfrac{N_{1}(A)\ n_{0}(B)}{n_{1}(A)}$
which is asymptotically normal with mean $\theta(B)=\mu_{0}(B)$ and standard error
$\text{SE}[\tilde{\theta}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+w_{1}\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$v_{0}=\dfrac{\theta(B)-M(B)}{M(B)}$
The estimator for the subregion total, $\tilde{\tau}(B)=\tilde{\theta}(B)+\tilde{N}_{1}(B)$, is also asymptotically normal with mean $\tau(B)$ and standard error
$\text{SE}[\tilde{\tau}(B)]=\theta(B)\sqrt{v_{0}\left(\frac{1}{p\,\theta(B)}\right)+\left(\frac{v_{1}\,N_{1}(B)\left(N_{1}(A\backslash B)-\theta(B)\right)^{2}+v_{2}\,N_{1}(A\backslash B)\left(N_{1}(B)+\theta(B)\right)^{2}}{N_{1}(A)\,\theta(B)^{2}}\right)\left(\frac{1-p}{p\,N_{1}(A)}\right)}$
where
$v_{1}=\dfrac{N_{1}(B)-M(B)}{M(B)+1}\qquad\text{and}\qquad v_{2}=\dfrac{N_{1}(A\backslash B)-M(A\backslash B)}{M(A\backslash B)+1}$
See Section 7.3.4 for details.
@@ 5. Discussion
We estimate 23,000 vendors work in New York City. In comparison, the 2017-2021 ACS estimates 4,634 individuals work the occupation of Door-To-Door Sales Workers, News And Street Vendors, And Related Workers in New York City. This suggests that, according to the ACS, there are at most 4,634 street vendors in New York City, and thus our finding indicates that the ACS estimate misses 80% of street vendors or more.
A fairer comparison might account for the fact that the standard error of our estimate is approximately 1,000, and the standard error of the ACS estimate is approximately 488, which we obtained from the replicate weights. Using the lower limit of a 95% confidence interval for our estimate (21,000) and the upper 95% confidence limit for the ACS estimate (5,610), our findings suggest the ACS misses roughly three quarters of street vendors or more.
Note the importance of the assumptions stated in Section 3.2 since other derivations of ratio estimation may imply smaller or larger standard errors, from which we might conclude the ACS misses a smaller or larger portion of the street vending population. For example, if the standard errors from Section 4.2 hold, then the confidence intervals would overlap if there were more than 300 vendors in the average market. We believe this is highly unlikely, suggesting the two estimates are inconsistent even under overdispersion.
Additional evidence comes from the OATH data, which indicate approximately 2,500 (unique) vendors are summoned to court each year for violating vending laws. If the ACS estimate holds, then more than half of vendors are summoned to court each year. If our estimate holds, then approximately one in ten vendors are summoned to court each year. We believe the latter is more realistic.
@@
Another explanation is that the Street Vendor Project survey and the ACS may be measuring different populations or concepts. For example, the ACS estimate reflects the average number of vendors between 2017 and 2021, while our estimate reflects 2020 and 2021. For another example, the Census Bureau determines the occupation of an ACS respondent by autocoding responses to write-in questions. It is possible that the Street Vendor Project’s definition of a vendor is more inclusive. Street vending often provides supplemental income, and thus the discrepancy may reflect the broader challenge of studying the gig economy. In this case, either estimate may be relevant, depending on the intended use case.